OpenStudy (anonymous):

Julie drives 100 mi to Grandmother's house. On the way to Grandmother's, Julie drives half the distance at 33.0 and half the distance at 80.0 . On her return trip, she drives half the time at 33.0 and half the time at 80.0 . What is Julie's average speed on the way to Grandmother's house? What is her average speed on the return trip?

OpenStudy (anonymous):

\[t=\frac{d}{r},r=\frac{d}{t} \]\[\frac{100}{\frac{50}{33}+\frac{50}{80}}=\frac{5280}{113}=46.7 \text{ mph}\]Let t be the total time for the return trip from Grandma's house:\[r t = d\]Solve the following for t:\[33 \frac{t}{2}+80\frac{t}{2}= 100\]\[t=\frac{200}{113} \]\[\frac{100}{\left(\frac{200}{113}\right)}=\frac{113}{2}=56.5 \text{mph} \]